Summary of the GMVA Session, Sep. 25-29, 2014 ============================================== Noto and Onsala not participating due to telescope maintance work, so the following stations observe in this session: Effelsberg, Yebes, Pico, PdBure, VLBA, GBT General Summary: ================ Fingetest reveals fringes to all European stations with SNR of up to 337, see table below. The dual polarisation recording at Pico now works again (using the MK5B recorder, the MK5C has a problem). Yebes has swapped polarisation, should be LCP but is RCP. Weather at Yb was ok. Pico had good weather on the first day, but after that stopped most of the time. PdB observed, except for MM001A (Marscher et al., Fermi sources), since the PC did not like the proposal. Weather at PdB was ok. GBT observed, see details below. VLBA observed, see details below. ---------------------------------- Daily notes: 25.9 PV, PdB, EB, PB, Yb start in time fringestest: fringes to all stations found Yb uses old SIS receiver, polarisation is swapped (should be LCP, but is RCP) PV uses MK5B+, both polarisations now work equally good SNR from Fringetest: EB-PICO RR SNR 337 t=30sec 3C454.3 EB-Yebes LL SNR 114 t=129sec 3C454.3 Pico-PdB LL SNR 230 t=1.2sec 3C454.3 26.9 8 UT 3C84 Pico good all night, Tsys 100K PdB good weather, Tsys 100K EB good weather, Tsys 200 K Yb good weather, Tsys 250 K, uses DSB tuning, must multiply Tsys from log by 2!! VLBA weather looks good, BR may have some humidity 16h00 UT Pico stop, bad weather 17h00 UT EB it started raining (CygA) 21h00 UT EB weather still bad FD, KP rain PT overcast NL, LA, OV , BR, MK ok 27.9 8 UT EB, rainy conditions, ptg.on OJ287/4C39.25 ok VLBA no rain, MK wind 11km/s 13 UT EB, clearing up, ptf and focus ok on OJ287/4C39.25, Tsys 230-240K FD overcast NL, LA, PT clear KP cloudy 17h30 UT PdB stopps for local project (they don't observer Marschers project), weather good 20 UT EB clear GBT joined 19h30 UT, weather good NL, FD, LA, PT, Mk, BR ok KP, OV clouds 28.9 8 UT EB last night and now clear and good PV resumes on 3C120, weather poor Yb observers, mixed weather, no obs. 0130-0320 UT on Sep 28, DSB Tsys 200K 1700 UT EB clear NL, FD, BR, MK clear OV, PT cloudy KP, LA overcast 2000 UT VLBA no rain 29.9 8 UT EB light clouds, last night was good, ok PV rejoin schedule, Tsys 140K, weather poor VLBA no rain, OV 0.02cm perhaps some 1130 UT EB, more cloudy, cannot see OJ287 1430 UT EB finished in cloudy weather Yb light rain in night, now clear sky (receiver no helium 07:30 - 09:30), Tsys 180-200 K (DSB) ------------------------------------- Individual station comments: Effelsberg: ---------- Sep 27: Effelsberg is observing, but the weather is not very good. Light rain and clouds during the night and now. An improvement is expected after noon. During the last night we could not see 2013+37, but now we can see and point on 4C39.25/OJ287. The system is working fine, Tsys = 220 K. Sep 28: Effelsberg observed last night in clear sky and excellent conditions, The weather is still very good, Tsys 300 K at elv=20 (3C120). Sep 29: EB stopped observing according to the schedule, during the day the cloud coverage increased and pointing become more difficult. But all in all the session was quite sucessfull for Effelsberg, this time. PdB: ----- Sep 25: Reference antenna pad is N07. We are on Plateau de Bure and the weather is ok (clear sky, phases ok). We had some power & meteo station problems in the afternoon that got resolved shortly before the session started. We are running the FS on the Mark5 now, so you won't have a GPS-PPS clock signal. Our current maser drift is low, however. We missed the first scan and the efficiency of scans 2 and 3 is not optimum, but from then on the data are ok. We started with 4 antennas but since 19:00 UT we are observing with 5 antennas. Sep 26: Plateau de Bure had good weather all night and observed with 5 antennas. We could not do one observation at high elevation, and we sacrificed the scan on 6:45 UT to redo a tuning of the receivers (one antenna had a Trec that was not optimum). Tsys is now 100K, clear sky, phases stable. Sep 27: Plateau de Bure observed until 17:22 UT local time under correct weather conditions (afternoon phases were turbulent - phasing the array was difficult at low elevation), then switched back to local projects as foreseen. We will join again on Monday. Pico: ----- Sep 26: We had good weather all the night and the observation+recording is running smoothly. Tsys about 100K. Sep27: Unfortunately the weather is bad here and we are stopped since yesterday 16:00. The forecast is not good for today. Sep 28: We just resumed the observation but in very poor weather conditions. 3c120 is still visible. Sep 29: We are joining again to the schedule. The weather is not good, with thick clouds most of the time. Tsys about 140K. Yebes: ------ Sep26: Yebes observed all night. Conditions were clear sky. Tsys ~ 240 K. Scan 63 lost. There has been a problem with the tsys computation up to 11:45 today. Almost all Tsys in our log file up to that hour are wrong. I do not know if I can rescale the wrong values and obtain correct ones. I have to think about it. From that hour onwards Tsys are correct. Sep 27-28: Yebes has observed on Saturday and Sunday, except between 1:30 UTC - 3:20 UTC today due to a hardware problem at the antenna. Weather conditions since Friday night are cloudy sky with clears and clouds and showers in some ocasions. Tsys ~ 200 K in DSB. SIS receiver is tuned to DSB. As for Tsys between 2014.268.18:00:13.03 and 2014.269.11:30:12.18, you will have to throw all numbers above 150, multiply by 2 and do an interpolation between those values which are "reasonable". Sep 29: Weather conditions in Yebes have changed from clouds with a short light rain during the night to clear sky now. The receiver has run out of helium and Tsys has risen between 07:30 and 09:30. The antenna was stopped between 09:10 and 09:30. After refilling the receiver with helium, Tsys has dropped again to 180-200 K (DSB). GBT: ---- Sep 27: The GMVA project C142A has started at Green Bank. We were unable to get a good AutoOOF solution, after a few trials, so are running with whatever surface we started with. We joined the observing schedule at 19:30 UT. So we can hope for better luck with the AutoOOFs for the sessions tonight. The weather is clear to partly cloudy with reasonably calm winds. Initial estimate of Tsys is about 130K. Sep 29: In general things went well at Green Bank. We had quite good weather for most of the run, except for the last hour or so this morning (Sept 29) when it started raining. The night Sept 27-28 was very clear, and the peak scans all looked good. During the day time, there were partly cloudy conditions, and some of the pointing scans were questionable. Some of the scans, and most of the calibration data, is missing from the portion on Sept 28, 14:00 to 16:00 UT due my error making a typo in the source name in editing of the script. The disk pack MPI+1621 is being shipped to Bonn today. VLBA: VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C142A Observer: Krichbaum, T. Project type: OTHER Obs filename: c142acrd.* Date/Day: 2014SEP26/269 Ants Scheduled: NL FD LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Betty Ragan. Vclock, gps delta, UT1-UTC all check OK. 0056 Begin at BR,PT,OV,MK 0056-0623 *OV 3mm cryo is cooling. (C135381) CRYO 50 163 0057 Begin at FD,KP,LA,NL 0112-0129 %BR raining. 0130 %OV, MK windy. 0340-0432 *PT,MK removed to run USNO project N4269. SCH 200 104 0449-0454 *OV DBE-1 had TimeStampSyncError alert. TIME 100 5 Resynced the timing. 0454-0520 *OV time error cleared, but tsys were showing TIME 100 26 as infinite, presumably due to timing-related issues. Temporarily stopping the observation on the control computer, and rebooting the DBE. 0452-0455 *KP had zero length recording scan. REC 100 3 0545 New operator is Alan Kerr 0546-0549 *PT autostow. power glitch. COM 100 3 0556-0557 *PT autostowed. Power glitch. COM 100 1 0623 OV 3mm has cooled. 0642-0657 *PT autostow. power glitch. COM 100 15 1254 %BR 7mm. SYN2 is not locking at 7.6 with the 100 MHz IF reporting 0.229. Site Tech says this is a known problem but will investigate. 1338 New operator: T. Perreault 1405 BR wea: broken low clouds, no precipitation 1425 LA wea: clear, no precipitation 1428 FD wea: overcast, damp, no current precipitation 1455 OV wea: cloudy, no precipitation 1516 KP wea: scattered low clouds, no precipitation 1559-1602 *OV autostowed, power glitch, repointed COM 100 3 1914-2038 *PT removed from observation so that site tech CRYO 100 84 can replace 4cm coldhead (C135387) 2007 KP wea: low overcast, rain and lightning 2115 %FD raining. 2118-2122 *KP autostowed, power glitch, repointed COM 100 4 2150 New operator is Betty Ragan. 2249-2302 *PT had DBE1-TimestampSyncError. Resynched the TIME 100 13 clocks, but the error did not go away, and the tsys were showing as infinite, as with OV yesterday. Rebooting the DBE. 2315 %KP raining. 0000 Date is 2014SEP27/270. 0100 %OV windy. 0545 New operator is Alan Kerr 0955-1049 *PT,MK out for scheduled USNO observation. SCH 200 108 1352 Operator is A. Sowinski. 1707-1710 *BR Mark5C python exception and zero record rate. REC 100 3 Restarted DRS, checked module (MPI+1618) for Smart errors, didn't find any. Error cleared. 1715-1721 *BR Same problem. Restarted DRS and switched to REC 100 6 NRAO+383 in bank A. Seems fine now. 1950 *KP Raining. 2014-2222 *KP Autostow, power glitch. Can't repoint - COM 100 128 seems to be a blown fuse in AZ#2 (C135393). 2204 MK Windy. 2220 KP Site tech has replaced the ACu fuse. 2223-2226 *KP ACU EL#1 armature/electronics fault. ACU 100 3 2256 KP Windy. 0000 Date is 2014SEP28/271. 0145 New operator is Alan Kerr 0225-0318 *PT,MK out for scheduled USNO observation. SCH 200 106 0635-0637 *MK autostow. AZ#1 and EL#2 armature/electronics CB ACU 100 2 tripped. reset, repointed. 1323 *KP Windy. 1336-1339 *MK autostow. power glitch. COM 100 3 1348 Operator is A. Sowinski. 1506-1510 *KP Autostow with AZ#1 and EL#1 CB faults. Possibly ACU 100 4 due to winds causing high motor currents. Faults cleared; repointed. 0000 Date is 2014SEP29/272. 0012 *OV light rain. 0052-0055 *BR Mark5C zero record rate for this scan. No REC 100 3 other errors or alerts. 0145 New operator is Alan Kerr 0310-0322 *BR zero-length-error. cleared by itself after two TIME 100 12 scans. 0355-0449 *PT,MK out for scheduled USNO observation. SCH 200 108 1130 End at GB 1340 New Operator: Paul Padilla 1459-1501 *FD autostowed, power glitch? COM 100 2 2033 %LA raining 2055 End. Downtime Summary: Total downtime : 912 min Percentage downtime of observing: 2.1% Average downtime per hour : 1.2 min Total scheduled observing time (# Antennas): 44152 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.